OpenStudy (anonymous):

Given the following identity, secxcscx(tanx+cotx) = 2 + tan^2(x) + cot^2(x): Prove the identity by completing the table below, indicating the steps on the left and the reasoning on the right.

OpenStudy (anonymous):

\[secxcscx(tanx+cotx) = 2 + \tan^2x + \cot^2x\]

OpenStudy (anonymous):

I'm exactly sure how I'm supposed to start...

OpenStudy (ankit042):

start by expanding the terms

OpenStudy (ankit042):

you will get secXcscXtanX +secXcscXcotX now convert everything in terms of sinX and cosX

OpenStudy (anonymous):

\[\csc^2(x)+\sec^2(x) \]this is what I got!

OpenStudy (anonymous):

\[\sec x \csc x \left( \tan x+\cot x \right)=\frac{ 1 }{\cos x \sin x }\left( \frac{ \sin x }{\cos x }+\frac{\cos x }{\sin x } \right)\] \[=\frac{ 1 }{\cos ^{2}x }+\frac{ 1 }{\sin ^{2} x}\] \[=\sec ^{2}x+\csc ^{2}x=1+\tan ^{2}x+1+\cot ^{2}x=R.H.S\]

OpenStudy (ankit042):

yes now you can solve the right hand side use 1+tan^2=sec^ and 1+csc^2X=csc^2X

OpenStudy (anonymous):

\[\csc^2(x) \sec^2(x) \] !!!

OpenStudy (anonymous):

But then which parts go in the boxes?